OpenStudy (faithspring):

Find the slope: 5. (4, -1) and (2, 6) 6. (0, -3) and (9, 0)

OpenStudy (trojanpoem):

\[\frac{ \Delta y }{ \Delta x } = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\]

OpenStudy (trojanpoem):

In your case 6-4/2-(-1) = 2/3 0 +3 / 9 - 0 = 1/3

OpenStudy (faithspring):

thank you, that helped a lot